# Graph Foundation Models

### v2

@@ 3.4 Transferability principle in graph classification
Stability. Huang et al. (2023d) proposes a provably stable position encoding that surpasses the expressive sign and invariant encoding (Kreuzer et al., 2021) and modeling (Lim et al., 2022), enabling minimal changes to positional encodings on the minor modifications to the Laplacian. The key innovation is to apply a weighted sum of eigenvectors instead of treating each eigensubspace independently. Satisfying performance can be observed on the out-of-distribution molecular graph prediction.
@@
3.5 The transferability across tasks
A unified task formulation is generally employed to facilitate transferability across various tasks. The significance of aligning task formulations is evident in the following example: Jin et al. (2020b) shows that using link prediction directly as a pretext task leads to negative transfer for node classification. However, by reformulating node classification into a link prediction problem (Sun et al., 2022; Huang et al., 2023a), where a node’s class membership is treated as the link likelihood between the node and label nodes, positive transfer is achieved.
Liu et al. (2023g); Sun et al. (2023) further propose a subgraph view to adapt the node classification as an ego graph classification, and link prediction as a binary classification on the induced subgraph of the target node pair.
Figure 2 provides illustrative examples for these two unified views. The unified task formulation enables (1) enlarging the dataset size via converting datasets for different downstream tasks as one and (2) utilizing one pre-training model to serve different tasks.
Surpassing the unified formulation, there are shared principles on different tasks listed as follows: (1) Node classification and link prediction tasks share the feature homophily as an important principle. (2) Liu et al. (2023d) indicates that the global structural proximity principle on the link prediction can improve the node classification performance on the non-homophilous graph. (3) The triadic closure in the link prediction is a particular network motif utilized in the graph classification. There are more shared motifs (Hibshman et al., 2021; Dong et al., 2017; AbuOda et al., 2020; Kriege et al., 2020) on both graph classification and link prediction tasks.
4 Towards GFM following neural scaling law
The success of the foundation model can be attributed to the validity of the neural scaling law (Kaplan et al., 2020) which shows performance enhancement with increasing model scale and data scale. In this section, we first discuss when the neural scaling law happens in Section 4.1 from a graph vocabulary perspective. We then discuss techniques towards successful data scaling and model scaling in Section 4.2 and 4.3, respectively. The pre-text task design is finally discussed in Section 4.4.
@@ 4.1 When neural scaling law happens
In section 3, we discuss the underlying transferable principles guiding future vocabulary construction. Such principle guidance has led to the successful scaling behavior in the material science domain (Shoghi et al., 2023; Zhang et al., 2023a; Batatia et al., 2023) with the help of the geometric prior. Nonetheless, we are still cautious about whether the existing success can be extended to the graph domain. The key concern is whether graphs can strictly follow those principles. Uncertainty can be found on the human-defined graph construction criteria (Brugere et al., 2018). For instance, the construction knowledge relying on expert knowledge may lead to uncertainty in edges (Ye et al., 2022). Chen et al. (2023b); Li et al. (2023c) observe that the mislabeled samples widely exist across datasets, where the popular CiteSeer dataset has more than 15 Despite the above uncertainty, different graph constructions with manual design can follow opposite principles. For instance, Ogbn-arxiv (Hu et al., 2020) and Arxiv-year (Lim et al., 2021) are two node classification datasets with identical graph information. The only difference lies in the label where Ogbn-arxiv employs paper categories, and Arxiv-year uses publication years as labels, resulting in conflicting homophily and heterophily properties (Mao et al., 2023a). Therefore, when uncertainties and opposite graph constructions exist, the scaling behavior may not happen as the data does not obey the graph transferable principles.
@@ 4.2 Data scaling
Chen et al. (2023a); Huang et al. (2023a) initially validate that GNNs trained in both supervised and self-supervised manners follow data scaling law on molecular property predictions, and node classification on text-attributed graphs. Xu et al. (2023) further exhibits that the similarity between pre-training data and downstream task data serves as a prerequisite for the data scaling on graphs. Specifically, Cao et al. (2023); Xu et al. (2023) provide concrete guidance on how to select the pre-training data via the graphon signal analysis and the essential network property, i.e. network entropy, respectively. The positive transfer can be found when the data exhibits similarity in the graphon space or with shared network properties.
A limitation in current research on data scaling is the small dataset size, particularly in contrast to the pre-training for LLMs (Raffel et al., 2019). This disparity may be attributed to the scarcity of high-quality data in the graph domain. NetworkRepository (Rossi & Ahmed, 2015), the largest graph database presently available, may support larger-scale pre-training, while burdensome pre-processing is required to clean those noisy and disordered data. Moreover, we illustrate a collection of useful datasets in Appendix A. We then show existing solutions as follows.
Tackling feature heterogeneity. Existing graph datasets cannot be uniformly utilized for pre-training due to the heterogeneous feature induced by missing features and different semantic spaces. Feature imputation techniques (Taguchi et al., 2021; Um et al., 2023; Gupta et al., 2023) are generally adapted to predict the missing attributes based on neighboring features. However, those techniques require each feature dimension to share the same semantic meaning. When features are not aligned in the same semantic space, Liu et al. (2023b) manually converts the original features with text descriptions and then encodes the embedding with LLMs. Moreover, feature misalignment can also be found in the inference stage between the pre-training model input and the test data. Jing et al. (2023) concatenates a learnable padding feature on the downstream task feature to align with the pre-trained GNN. However, such a technique cannot adapt to the case when the feature space is not aligned. Zhao et al. (2023b) directly abandons the original feature and utilizes the feature similarity as guidance.
Generating synthetic graphs is another path to enhance data scaling and enlarge the training data size. Traditional graph generative models (Albert & Barabási, 2002; Robins et al., 2007; Airoldi et al., 2008; Leskovec et al., 2010) are capable of generating graphs satisfying some certain statistical properties, which still plays an important role on node-level and link-level tasks.
Deep generative models on graph (Jin et al., 2020a; Luo et al., 2021; Jo et al., 2022; Vignac et al., 2023) have shown great success in generating high-quality synthetic graphs which helps graph-level tasks by providing a more comprehensive description of the graph distributions space.
Liu et al. (2023a) utilizes diffusion models to generate task-specific data argumentation to enhance the GNN on graph-level tasks with significant performance gain. With successful evidence of pre-training on synthetic data from other domains (Mishra et al., 2022; Trinh et al., 2024), we anticipate the potential to generate high-quality synthetic graphs for large-scale pre-training towards GFM.
@@ 4.3 Model scaling
Previous research in NLP indicates that apart from data, the backbone model constitutes a fundamental for scaling (Kaplan et al., 2020). Liu et al. (2024) primarily validates the neural scaling law on various graph tasks and model architectures under the supervised setting.
@@
However, Kim et al. (2022) demonstrates that the GAT (Veličković et al., 2017) with a larger number of parameters underperforms on the graph regression tasks compared to the smaller-sized counterparts. As a comparison, geometric GNNs scale well to predict atomic potentials in material science (Shoghi et al., 2023; Zhang et al., 2023a; Batatia et al., 2023). The contradictory phenomenon indicates that the alignment between the geometric prior of GNNs and the specific task may influence model scaling.
@@
Graph transformer is another popular choice for the model architecture, where geometric priors are explicitly modeled through either GNN or positional encoding (Müller et al., 2023). Masters et al. (2022); Lu et al. (2023) show that graph transformers show positive scaling capabilities for molecular data under a supervised setting. More recently, Zhao et al. (2023c) demonstrates vanilla transformer’s effectiveness in protein and molecular property prediction. Particularly, it views the graph as a sequence of tokens forming an Eulerian path (Edmonds & Johnson, 1973), which ensures the lossless serialization, and then adopts next-token prediction to pre-train transformers. After fine-tuning, it achieves promising results on the protein association prediction and molecular property prediction and shows that vanilla transformers also follow the model scaling law (Kaplan et al., 2020). Nonetheless, the effectiveness of such pure transformers on other tasks remains unclear.
4.4 Pretext task design
The generative self-supervised learning designs the pre-text tasks (Hou et al., 2022; Hu et al., 2019; Kipf & Welling, 2016) to capture the shared data generation process among different tasks. Particularly, they attempt to predict the masked portions of the graph using the remaining structure and features. Liu et al. (2023f); Xia et al. (2023) further observe that task granularity also plays an important role in generative modeling. Specifically, employing node-level pretext tasks may lead the model to learn only low-level features (Liu et al., 2023f) while ignoring the global information essential for graph-level tasks. To address this issue, they adopt a GNN-based tokenizer to explicitly model high-level information in the pre-training stage and thus improve the downstream task performance.
More recently, the next token prediction (NTP) pre-text task (Zhao et al., 2023c) achieves initial success in the molecular graph. Notably, this is the first pretext task demonstrating empirical evidence of model scaling. The potential reason for its success may be (1) the construction of a fixed token set, narrowing down the problem space in a finite set to only predict a discrete token and (2) choosing transformers as the backbone model. However, it remains unclear whether the success can be easily extended to more tasks.
@@ 5 Insights & open questions
The usage of LLM in the graph domain. With the existing success of LLMs (Bubeck et al., 2023) in the NLP domain, a natural question is whether we can directly apply its original language-centric ability on the graph. Primary exploration of the LLM focuses on text-attributed graphs. Chen et al. (2023b); He et al. (2023) treat node classification as the text classification on the target node, illustrating promising results on the zero-shot setting even without any structural information. For graphs without text attributes, Fatemi et al. (2023); Wang et al. (2023a) feed the data into LLM via describing graph structure with natural language. Observations show that LLMs present limited capability in understanding the essential graph structures. The key reason is that the LLM is required to tackle problems sequentially greedily (McCoy et al., 2023), leading to a shortcut solution (Saparov & He, 2022; Dziri et al., 2023) rather than a formal analysis on the graph structure. Moreover, Yue et al. (2023) points out the efficiency issue of utilizing LLM on the recommendation, the downstream link prediction task.
Tang et al. (2023); Ye et al. (2023); Chai et al. (2023) thus utilizes the GNN to encoder the graph structure knowledge and then introduces a linear layer to transform the embeddings obtained from the GNN into text space as a prompt token. Although these models can perform well, they still have two shortcomings: (1) The ability to process structures is bounded by the capabilities of GNN; (2) The instruction tuning can be costly while the tuned model can only tackle the corresponding downstream task and is not transferable to other tasks and datasets, which makes their capabilities distant from a GFM. We agree that LLM illustrates superior performance on textual node feature understanding. Nonetheless, it remains unclear whether LLM should play a key role in building GFM or just serve as a better textual feature encoder.
A potential strategy for equipping LLMs to effectively handle graph structures could be inspired by the success of multi-modal foundation models (Liu et al., 2023c). The key idea (Yu et al., 2023b) is to map the graph structure to a series of discrete tokens with a pre-trained tokenizer, which may convert graphs into a format comprehensible by LLMs. More discussion can be found in Appendix C.
Efficiency issue in subgraph-based methods. Subgraph-based extraction is a widely adopted technique to achieve inductive inference (Zeng et al., 2021) and unify different task formulations (Sun et al., 2023). Nonetheless, the subgraph-based extraction leads to the following issues: (1) information loss in high-order neighborhoods; (2) duplicate subgraph information with excessive memory consumption, and (3) the time complexity of vanilla subgraph extraction grows exponentially with the number of hops. Those issues hinder the applicability of subgraph-based methods. Graph sampling techniques like (Zeng et al., 2019) and global state vectors (Fey et al., 2021) can help to alleviate these issues.
@@
Potential redundancy on pre-text task and architecture design. There are mainly two approaches to achieving transferability: (1) designing GNNs with specific geometric properties for transfer, e.g., ULTRA (Galkin et al., 2023), and (2) creating pretext tasks to automatically learn these properties. (Jin et al., 2020b) suggests an overlap between these approaches, indicating that pretext tasks targeting local structural information might be unnecessary, given that GNNs often inherently encode this information. Investigating the strengths and limitations of these techniques, along with providing practical guidance for their selection, could be a valuable research direction. A hypothesis might be that model design methods are more suitable for data that strictly adheres to geometric priors, while pretext task designs are more effective in the opposite scenario.
@@
The necessity of GFM. Graphs can be defined in different ways based on different criteria like similarity or influence between node pairs (Brugere et al., 2018). We can then categorize graphs based on the observability of the criteria. The observable graph is unambiguously known, e.g., whether one paper cites another paper in a citation graph. Text and images can also be viewed as a specific case of observable graphs. In contrast, the unobservable ones are manually conducted with ambiguous descriptions of the relationship, e.g., whether one gene regulates the expression of another in a gene-regulate graph. These graphs may not naturally exist in the world, leading to uncertainty with a lack of invariant principle. We are not sure whether such graphs should be considered in the GFM.
@@
Moreover, there are concerns about the benefit of training a GFM on graphs that are neither from the same domain nor share the same downstream task. On the one hand, it seems that training on them simultaneously shows no positive transfer benefit while increasing the risk of the negative transfer. On the other hand, we are not aware of potential undiscovered transferability leading to success.
GFMs on more applications. In this paper, we majorly focus on the conventional graph tasks, e.g., node classification, link prediction, graph classification, and the principle for building GFM on them. However, graph formulation is the universal representation ability, which can easily extend to different domains, e.g., scene graph for CV (Zhai et al., 2023; Zhong et al., 2021), bipartite graph for linear programming (Chen et al., 2022), and physical graphs for understanding physical mechanisms (Shi et al., 2022). Each domain can have its principle requiring specific graph vocabulary construction. Existing success can be found in material science (Zhang et al., 2023a; Batatia et al., 2023). For instance, (Batatia et al., 2023) utilizes the Equivariant Graph Tensor Network which can encode many-body information of atomic geometry can conduct molecular dynamics simulation across different chemistries. A potential application of GFM is to aid the generation of scene graphs (Zhong et al., 2021), involving predicting relationships between localized objects. Knowledge graph foundation models (Galkin et al., 2023) which can generalize to unseen entities and relations, may effectively model relationships and thus generate high-quality scene graphs.

### v3

@@ 3.4 Transferability Principles in Graph Classification
Stability. Huang et al. (2023d) proposes a provably stable position encoding that surpasses the expressive sign and invariant encoding (Kreuzer et al., 2021) and modeling (Lim et al., 2022), enabling minimal changes to positional encodings on the minor modifications to the Laplacian. The key innovation is to apply a weighted sum of eigenvectors instead of treating each eigensubspace independently. Satisfying performance can be observed on the out-of-distribution molecular graph prediction. Such stable positional encoding may be directly applied towards better GFMs.
Actionable step inspired by principles. Inspired by the network analysis with graph kernels, one concrete next step towards GFMs could be revolving on how to identify frequent network motif (Hočevar & Demšar, 2014; Ribeiro et al., 2021) which should be transferable across all graphs. Expressive GNNs with better network motif model capability could be a suitable architecture towards GFMs.
@@
3.5 Transferability Principles across Tasks
A unified task formulation is generally employed to facilitate transferability across various tasks. The unified task formulation enables (1) enlarging the dataset size via converting datasets for different downstream tasks as one and (2) utilizing one pre-training model to serve different tasks. The significance of aligning task formulations is evident in the following example: Jin et al. (2020b) shows that using link prediction directly as a pretext task leads to negative transfer for node classification. However, by reformulating node classification into a link prediction problem (Sun et al., 2022; Huang et al., 2023a), where a node’s class membership is treated as the link likelihood between the node and label nodes, positive transfer is achieved. Liu et al. (2023f); Sun et al. (2023) further propose a sub-graph view to adapt the node classification as an ego graph classification, and link prediction as a binary classification on the induced sub-graph of the target node pair. Figure 2 provides illustrative examples for these two unified views. More recently, Liu et al. (2023b) unifies node-level, link-level, and graph-level tasks via (1) adding a virtual prompt node and (2) connecting the virtual nodes to nodes of interests, i.e., the center node for node classification, source and target nodes for link prediction, and all nodes for graph classification.
A unified formulation provides the possibility for co-training all tasks together while it remains unknown whether this can be done without negative transfer. Moreover, the unified task formulation may not be necessary to achieve transfer across tasks. It is generally utilized for supervised co-training and prompt-based prediction as discussed above. A GFM can be (1) pre-trained with self-supervised tasks and (2) adapted to downstream tasks via fine-tuning without requiring specific task formulations. The success is due to the transferability principles across different tasks. However, there remains limited study in this direction. We list a few existing principles as follows. (1) Node classification and link prediction tasks share the feature homophily as an important principle. (2) Liu et al. (2023c) indicates that the global structural proximity principle on the link prediction can improve the node classification performance on the non-homophilous graph. (3) The triadic closure in the link prediction is a particular network motif utilized in the graph classification. There are more shared motifs (Hibshman et al., 2021; Dong et al., 2017; AbuOda et al., 2020; Kriege et al., 2020) on both graph classification and link prediction tasks. We emphasize the importance of cross-task transferability principles as an important future direction.
4 Neural Scaling Law on GFMs
The success of the foundation model can be attributed to the validity of the neural scaling law (Kaplan et al., 2020) which shows performance enhancement with increasing model scale and data scale. In this section, we first discuss when the neural scaling law happens in Section 4.1 from a graph vocabulary perspective. We then discuss techniques towards successful data scaling and model scaling in Section 4.2 and 4.3, respectively. We finally discuss the potential on leveraging large-scale LM on the graph domain in Section 4.4. More discussions on technical details can be found in Appendix C.
@@ 4.1 When Neural Scaling Law Happens
In section 3, we discuss the underlying transferable principles guiding future vocabulary construction. Such principle guidance has led to the successful scaling behavior in the material science domain (Shoghi et al., 2023; Zhang et al., 2023a; Batatia et al., 2023) with the help of the geometric prior. Nonetheless, we are still cautious about whether the existing success can be extended to the graph domain. The key concern is whether graphs can strictly follow those principles. Uncertainty can be found on the human-defined graph construction criteria (Brugere et al., 2018). For instance, the construction knowledge relying on expert knowledge may lead to uncertainty in edges (Ye et al., 2022). Chen et al. (2023b); Li et al. (2023c) observe that the mislabeled samples widely exist across datasets, where the popular CiteSeer dataset has more than 15% wrongly labeled data. Despite the above uncertainty, different graph constructions with manual design can follow opposite principles. For instance, Ogbn-arxiv (Hu et al., 2020) and Arxiv-year (Lim et al., 2021) are two node classification datasets with identical graph information. The only difference lies in the label where Ogbn-arxiv employs paper categories, and Arxiv-year uses publication years as labels, resulting in conflicting homophily and heterophily properties (Mao et al., 2023a). Therefore, when uncertainties and opposite graph constructions exist, the scaling behavior may not happen as the data does not obey the graph transferable principles.
@@ 4.2 Data Scaling
Data scaling refers to the phenononmon that the performance consistently improves with the increasing data scale. Chen et al. (2023a); Huang et al. (2023a) initially validate that GNNs trained in both supervised and self-supervised manners follow data scaling law on molecular property predictions, and node classification on text-attributed graphs. Cao et al. (2023) further exhibits that the similarity between pre-training data and downstream task data serves as a prerequisite for the data scaling on graphs. Specifically, Cao et al. (2023) provides concrete guidance on how to select the pre-training data via the graphon signal analysis and the essential network property, i.e. network entropy, respectively. Notably, all principles mentioned in Section 3 can be applied to facilitate positive transfer with data scaling phenomena.
A limitation in current research on data scaling is graph data insufficiency, in contrast to the readily available trillion-level real-world data in CV and NLP domains. The key reasons are two-fold: (1) constructing graphs requires expert intervention e.g., defining relationships (2) intellectual property issues. We endeavour to collect all the open-source graph datasets with details in Appendix A.
Synthetic graph generation can be utilized to alleviate the data insufficiency issue, enableing more comprehensive training. Traditional graph generative models (Albert & Barabási, 2002; Robins et al., 2007; Airoldi et al., 2008; Leskovec et al., 2010) are capable of generating graphs satisfying some certain statistical properties, which still plays an important role on node-level and link-level tasks. Deep generative models on graph (Jin et al., 2020a; Luo et al., 2021; Jo et al., 2022; Vignac et al., 2023; Liu et al., 2023a) have shown great success in generating high-quality synthetic graphs which helps graph-level tasks by providing a more comprehensive description of the graph distributions space. With successful evidence of pre-training on synthetic data from other domains (Mishra et al., 2022; Trinh et al., 2024), we anticipate the potential on benefits from high-quality synthetic graphs.
@@ 4.3 Model Scaling
Model scaling refers to the phenomenon that the performance consistently improves with the increasing model scale. Previous research in NLP indicates that apart from data, the backbone model constitutes a fundamental for scaling (Kaplan et al., 2020). Liu et al. (2024a) primarily validates the neural scaling law on various graph tasks and model architectures under the supervised setting.
@@
However, Kim et al. (2022) demonstrates that the GAT (Veličković et al., 2017) with a larger number of parameters underperforms on the graph regression tasks compared to the smaller-sized counterparts. As a comparison, geometric GNNs scale well to predict atomic potentials in material science (Shoghi et al., 2023; Zhang et al., 2023a; Batatia et al., 2023). Observations indicate that geometric GNNs with a good geometric-prior vocabulary design can help achieve model scaling over the vanilla GNN.
@@
Graph transformer is another popular choice for the model architecture, where geometric-prior graph vocabulary design is explicitly modeled through either a GNN encoder or positional encoding (Müller et al., 2023). Masters et al. (2022); Lu et al. (2023) show that graph transformers show positive scaling capabilities for molecular data under a supervised setting. More recently, Zhao et al. (2023b) demonstrates vanilla transformer’s effectiveness in protein and molecular property prediction. Particularly, it views the graph as a sequence of tokens forming an Eulerian path (Edmonds & Johnson, 1973), which ensures the lossless serialization, and then adopts next-token prediction to pre-train transformers. After fine-tuning, it achieves promising results on the protein association prediction and molecular property prediction and shows that vanilla transformers also follow the model scaling law (Kaplan et al., 2020). Nonetheless, the effectiveness of transformers on other tasks remains unclear.
4.4 Leveraging Large-scale LMs for Graphs
LLMs with successful scaling behavior have achieved tremendous success in the NLP domain. Surprisingly, well-trained LLMs can be applied to other domains with satisfying performance such as time series forecast (Gruver et al., 2024a) and material science (Gruver et al., 2024b). Larger-scale LLMs can even capture key symmetries of crystal structures, suggesting that LLMs may posses a strong simplicity bias (Panwar et al., 2023) across domains by implementing Bayesian model averaging algorithm (Zhang et al., 2023c).
A recent line of research on GFMs focuses on leveraging strong capabilities of LLMs on graph tasks. Our discussions can be roughly categorized on LLM applications (i) conventional graph tasks (such as node, edge, and graph classification), and (ii) language-driven tasks like Graph Question Answer (GQA).
LLMs on conventional graph tasks. One natural way to utilize LLMs is as textual feature encoders (Chen et al., 2023b). Despite original node features may not be text, Liu et al. (2023b) manually converts them into knowledge-enhanced text descriptions and then encodes features into textual embedding. This LLM embedding approach offers the following benefits. (i) High feature quality helps achieve satisfying performance with vanilla GCN (Chen et al., 2023b). (ii) LLMs encode diverse original features into an aligned feature space, enabling training and inference across graphs from different domains without the feature heterogeneity problem. Notably, when LLMs are utilized as feature encoders for textural understanding, the scaling law does not happen (BehnamGhader et al., 2024), meaning that a larger model does not necessarily lead to better performance.
Another approach is to utilize LLMs as predictors which first fine-tunes LLM and then generate predictions in a natural language form. Chen et al. (2023b); He et al. (2023) treats node classification as text classification on the target node feature, illustrating promising results in the zero-shot setting. However, simply flattening graph structures into prompts does not yield additional improvement, remaining a large performance gap compared to well-trained GNNs (Chen et al., 2023b). To better encode the graph structure knowledge, methods such as GNN (Tang et al., 2023), graph transformer (Chai et al., 2023), and non-parametric aggregation (Chen et al., 2024b) are utilized as structure encoders. The encoded structural embeddings are then linearly mapped into text space as prompt tokens. LLMs generate predictions based on a concatenation of the prompt token and the textual instruction. Instead of additional graph modeling, Zhao et al. (2023a) employs a novel tree-based prompt design that transforms the graph into sequence while retaining important structural semantics. This approach indicates the potential for LLMs to understand particular graph structures. Overall speaking, a proper LLM fine-tuning can achieving satisfying graph performance while the efficiency may be a potential issue.
LLMs on language-driven graph tasks. Instead of adapting LLM for conventional graph tasks, LLMs can also be applied to language-driven tasks they originally skilled in, for example, Graph Question Answer (GQA). Fatemi et al. (2023); Wang et al. (2023a) apply LLMs on various GQA tasks, e.g, cycle check, and maximum flow, by describing graph structure with natural language. More recently, Perozzi et al. (2024) incorperates an external GNN tokenizer to encode graph information, achieving satisfying out-of-domain generalization to unseen graph tasks. Interestingly, Perozzi et al. (2024) illustrates that equivariance is not necessary when equipped with LLMs. (He et al., 2024) proposes new real-world challenging GQA tasks and a corresponding LLM-based conversational framework. This framework integrates GNNs and retrieval-augmented generation (RAG) to improve graph understanding and mitigate issues like hallucination, demonstrating effectiveness across multiple domains. Until now, most GQA challenges have focused on the abstract graphs without concrete descriptions for each node, creating to obstacle to o leveraging the extensive internal knowledge in LLMs. We call for more real-world GQA challengesm enabling better leverage LLM capabilities.
Despite the above successes, it remains concerns on the LLM’s capability on understanding the essential graph structures. Saparov & He (2022); Dziri et al. (2023) theoretically observe that the LLM is required to tackle problems sequentially greedily (McCoy et al., 2023), leading to a shortcut solution rather than a formal analysis on the graph structure. A more comprehensive discussion can be found in Appendix E.1.
@@ 5 Insights & Open Questions
In this section, we explore key insights gained from recent advancements in GFMs and highlight open questions that remain to be addressed in this evolving field. More comprehensive discussions can be found in Appendix E
5.1 Potential Redundancy on Pretext Task and Architecture Design
@@
There are mainly two approaches to achieving transferability: (1) designing GNNs with specific geometric properties for transfer, e.g., ULTRA (Galkin et al., 2023), and (2) creating pretext tasks to automatically learn these properties. (Jin et al., 2020b) suggests an overlap between these approaches, indicating that pretext tasks targeting local structural information might be unnecessary, given that GNNs often inherently encode this information. Investigating the strengths and limitations of these techniques, along with providing practical guidance for their selection, could be a valuable research direction. A hypothesis might be that model design methods are more suitable for data that strictly adheres to geometric priors, while pretext task designs are more effective in the opposite scenario.
5.2 The Feasibility of GFMs
@@
Graphs can be defined in different ways based on different criteria like similarity or influence between node pairs (Brugere et al., 2018). We can then categorize graphs based on the observability of the criteria. The observable graph is unambiguously known, e.g., whether one paper cites another paper in a citation graph. Text and images can also be viewed as a specific case of observable graphs. In contrast, the unobservable ones are manually conducted with ambiguous descriptions of the relationship, e.g., whether one gene regulates the expression of another in a gene-regulate graph. These graphs may not naturally exist in the world, leading to uncertainty with a lack of invariant principle. It remains unknown whether GFMs can learn shared knowledge while avoiding manually introduced noisy patterns.
@@
There are concerns about the benefit of training a GFM on graphs that are neither from the same domain nor share the same downstream task. On the one hand, it seems that training on them simultaneously shows no positive transfer benefit while increasing the risk of the negative transfer. On the other hand, there may be potential undiscovered transferable patterns that could lead to success. Therefore, we pose an open question whether there exists a universal structural representation space that can benefit all the graph tasks?
5.3 Broader Usage of GFM
In this paper, we majorly focus on building GFM for conventional graph-focused tasks. Notably, graph formulation provides the universal representation ability, which has a broader usage in other domains, e.g., scene graphs for Computer Vision (CV) (Zhai et al., 2023; Zhong et al., 2021), bipartite graphs for linear programming (Chen et al., 2022), and physical graphs for understanding physical mechanisms (Shi et al., 2022). To emphasize more broader usage of GFMs, we illustrate the potential advantaged usage of GFMs over existing foundation models in reasoning, computer vision, and code intelligence domains domains. Details can be found as follows.
Reasoning. Ibarz et al. (2022) proposes a task-specific GFM, focusing on neural algorithmic reasoning tasks. A strong reasoning capability can be found with effectiveness across sorting, searching, and dynamic programming tasks. We argue that this GFM following the theory of algorithmic alignment (Xu et al., 2020) may achieve better reasoning capability than the LLM merely relying on the textual inputs via retrieving concepts co-occur frequently in training data (Prystawski & Goodman, 2023).
Computer Vision. Scene graph is a data structure representing objects, their attributes, and the relationships between them within an image, facilitating CV tasks such as image understanding and visual reasoning. However, current research remains a naive scene graph modeling with vanilla GNNs with more emphasis on image modeling. We argue that the GFM on the scene graph may help to preserve global and local scene-object relationships (Zhai et al., 2024), avoiding the potential conflict or redundancy between multiple objectives which frequently appears on the recent popular Sora model (Brooks et al., 2024).
Code Intelligence. Graphs, e.g., code property graph (Liu et al., 2024b), control flow graph, and program dependency graph, play an important role in code-relevant tasks, e.g., vulnerability detection (Liu et al., 2024b), fault localization (Rafi et al., 2024), and code search (Ling et al., 2021). Compared to sequence-based modeling with LLMs, graphs can provide a complementary perspective on the overlooked essential code attributes such as syntax, control flow, and data dependencies. However, the graph modeling remains naive with unknown transferability across different program languages.
Overall speaking, GFMs demonstrate unique value compared to foundation models in other domains. However, they are limited to applications involving graph structure data. An exciting future topic is how to adaptively combine GFM with other foundation models across different modality towards a powerful Artificial General Intelligence (AGI).